# Surrogate-Assisted Targeted Learning for Nested Bridge Functionals under Administrative Censoring

### 15 Mar 2026 v2

@@ 8 Discussion
The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts. Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term that does not vanish faster than the leading sandwich term under the present nuisance specification. This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), resolves this gap by capturing the fluctuation-step variance the sandwich misses. Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime. At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this is a well-understood property of leave-one-out procedures in small samples and represents a safe failure mode. We recommend the cluster jackknife as the default inference procedure for SA-TMLE, with the sandwich as a fast approximation when $J\geq 50$.
@@
The surrogate-bridge framework applies wherever a delayed primary outcome is observed sparsely and an intermediate variable is broadly available; extension to other longitudinal designs beyond SW-CRTs is straightforward in principle. Two open directions are: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
@@
The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, and the DML cross-product remainder (Proposition 1 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
All analyses were conducted in R (version 4.3 or later). The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE, available at https://github.com/[author]/swcrtSurrTMLE. The package depends on SuperLearner (Polley and van der Laan, 2010) for ensemble estimation, lme4 for the GLMM comparator, and tmle for reference TMLE implementations. A fully reproducible script that replicates all tables and figures in this paper is included in the package vignette and archived on Zenodo at https://doi.org/10.5281/zenodo.[XXXXX] (DOI to be assigned upon acceptance).

### 31 Mar 2026 v4

@@ 8 Discussion
The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts. Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term of the same order as the leading sandwich term under the present nuisance specification. This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), substantially reduces this gap by capturing fluctuation-step variability omitted by the sandwich approximation. Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime. At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this pattern is consistent with the finite-sample conservatism often seen for leave-one-out procedures when $J$ is small and represents a safe failure mode. When accurate finite-sample coverage is a priority, we recommend the cluster jackknife for SA-TMLE; the sandwich remains a useful first-order approximation in larger-$J$ settings.
@@
Beyond stepped-wedge trials.
The nested bridge functional (1) and the SA-TMLE are not tied to the SW-CRT design. The identification conditions (Assumptions 15) and the $R_{SY}$ obstruction arise whenever three features co-occur: (i) a delayed primary outcome $Y$ is administratively censored for a subset of units; (ii) a shorter-term surrogate $S$ is observed broadly; and (iii) censoring is conditionally independent of $Y$ given $S$ (surrogate-mediated MAR). We briefly sketch two additional settings where these conditions hold.
Staggered-enrollment RCTs with administrative database linkage. Large pragmatic trials increasingly randomize individuals over a multi-year enrollment window and ascertain long-term outcomes (e.g., cardiovascular events, disability) through administrative claims linkage at a fixed analysis date. Participants enrolled in the final cohort have insufficient follow-up for the primary endpoint, generating design-induced administrative censoring analogous to the SW-CRT near-boundary regime. A short-term clinical measure collected at a scheduled visit (e.g., 6-month biomarker panel, early functional status) serves as the surrogate $S$. Because censoring is driven by enrollment date relative to the database lock — not by the participant’s health trajectory — Assumption 3 is as defensible as in the SW-CRT setting, the cluster index $j$ maps to enrollment cohort or site, and the SA-TMLE applies without modification. The $R_{SY}$ obstruction persists in this setting because the target functional retains the nested integral structure: the ATE is identified by integrating $\bar{Q}_{Y}(S,a,W)$ over the treatment-arm-specific distribution of $S$, not by inverse-weighting the censored observations.
Multi-site registry studies with delayed outcome ascertainment. Cancer registries and transplant databases routinely collect short-term post-treatment response indicators (e.g., pathological response, graft function at 3 months) for all patients, while long-term survival or recurrence data accrue differentially across sites depending on registry maturity and loss to follow-up. When follow-up incompleteness is driven by administrative reporting lags rather than informative dropout, the surrogate-mediated MAR condition holds with sites as clusters. The nested bridge functional identifies the treatment effect on the long-term outcome by integrating the outcome regression over the conditional distribution of the early response indicator, and the SA-TMLE provides doubly robust inference without requiring stable inverse-censoring weights across sites with heterogeneous follow-up completeness.
In both settings, the theoretical results of Sections 35 apply directly: the vanishing $\mathcal{T}_{\Delta}$ component (Lemma 1), the $R_{SY}$ obstruction (Proposition 2), the two-stage fluctuation resolution (Theorem 2), and the jackknife variance decomposition (Theorem 3) all follow from the functional structure, not from the specific SW-CRT data-generating process.
@@
Two open theoretical directions remain: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring, where the surrogate-mediated MAR condition must be reformulated in terms of a counting-process filtration; and (ii) providing a closed-form characterization of the fluctuation-step variance $V_{\varepsilon}$ in Theorem 3 as a function of the DGP and nuisance-estimation rates, which would allow analytic comparison of jackknife and sandwich efficiency without simulation. A third practical direction is establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
@@
The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, the DML cross-product remainder (Proposition 1 proof), and jackknife variance consistency (Theorem 3 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
All analyses were conducted in R (version 4.3 or later). The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE. The package depends on SuperLearner (Polley and van der Laan, 2010) for ensemble estimation, lme4 for the GLMM comparator, and tmle for reference TMLE implementations. Code and data sufficient to reproduce all tables and figures in this paper will be made available in a public repository upon acceptance. All simulation code is included as supplementary material with this submission.